# Discrete time crystals in the presence of non-Markovian dynamics

### v1

@@ II.2.3 Non-Markovian regime with noisy aperiodic $\kappa(t)$
We next focus on the question whether the DTC phase in the non-Markovian regime is stable in the presence of random noise in $\kappa(t)$. To this end, we consider an aperiodic $\kappa(t)$, realized by introducing random fluctuations in the dissipator coefficient, given by $\kappa(t)^{\prime}=\kappa(t)+a_{0}f(t)$.
@@
Figure 4: Figure showing (a) an aperiodic $\kappa(t)$, formed through the introduction of random fluctuations in the dissipation (see Eq. (5)), and (b) the corresponding DTC phase. Here $\kappa_{0}=2.7,m=\kappa_{0}/4,\epsilon=0.03$,$a_{0}=0.5$ and $\kappa_{max}=5$
@@ III Conclusion
We study the emergence of DTC in the presence of non-Markovian dynamics by considering an open Dicke model with Jaynes-Cummings-like dissipation $\kappa(t)$ (Cf. Eq. (3)). Our analysis shows that in contrast to Markovian dynamics, the DTC phase is more robust to a wide range of parameter values in the non-Markovian regime ($m<2\kappa_{0}$), which may be attributed to periodically varying $\kappa(t)$ even at long times. We present a dynamical phase diagram w.r.t. $\epsilon$ and $m$ (Cf. Fig. 3). The transition from Markovian to non-Markovian regime at $m=2\kappa_{0}$ is marked by TISS for all values of $\epsilon$ in the Markovian regime for large enough $\kappa_{0}$, which changes to varied dynamical phases, including DTC, in the non-Markovian regime. Furthermore, the DTC phase emerging for $m$ close to $2\kappa_{0}$ shows substantial robustness w.r.t. $\epsilon$, while this robustness is reduced, even though does not vanish, as we move deeper into the non-Markovian regime of $m<2\kappa_{0}$. Furthermore, we verify the presence of robust DTC order in the non-Markovian regime even for aperiodic $\kappa(t)$, realized by introducing a random noise in the dissipator coefficient.
@@
We expect the dissipative DTCs studied here can be realized experimentally in currently existing setups. For example, the presence of DTCs has been verified experimentally in different platforms, both in closed quantum systems, such as in ion traps [7], and in the presence of dissipation, for example in optical cavities [41]. Experimental studies of DTCs in the presence of different forms of dissipation would require control over bath spectral functions, which for example can be achieved through the introduction of filters [42].
Our results show that non-Markovian dynamics can be highly relevant for time-translational symmetry breaking in many-body open quantum systems, and can show features distinct from those seen in the Markovian regime. This also raises open questions regarding the fate of DTCs in more generic forms of non-Markovian dynamics, and regarding the possible role played by information backflow in the formation of DTCs.
@@ Appendix A Evolution equations and steady states
Appendix B Dynamics for large $\kappa_{max}$
In case of large $\kappa_{max}$ dissipation dominates the dynamics, such that the system reaches a TISS, as shown in Fig.5 a) or a thermal phase, as shown in Fig. 5 b).
Figure 5: Stroboscopic dynamics for $\kappa_{max}=10$, $\epsilon=0.02$ , a) $m=\kappa_{0}$ and b) $m=\frac{\kappa_{0}}{4}$.

### v2

@@ II.2.3 Non-Markovian regime with noisy aperiodic $\kappa(t)$
We next focus on the question whether the DTC phase in the non-Markovian regime is stable in the presence of random noise in $\kappa(t)$. To this end, we consider an aperiodic $\kappa(t)$, realized by introducing random fluctuations in the dissipator coefficient, given by:
@@
Figure 4: Figure showing (a) an aperiodic $\kappa(t)$, formed through the introduction of random fluctuations in the dissipation (see Eq. (6)), and (b) the corresponding DTC phase. Here $\kappa_{0}=2.7,m=\kappa_{0}/4,\epsilon=0.03$,$a_{0}=0.5$ and $\kappa_{\rm{max}}=5$.
@@ III Conclusion
We study the emergence of DTC in the presence of non-Markovian dynamics by considering an open Dicke model with Jaynes-Cummings-like dissipation $\kappa(t)$ (Cf. Eq. (3)). Our analysis shows that in contrast to Markovian dynamics, the DTC phase is more robust to a wide range of parameter values in the non-Markovian regime ($m<2\kappa_{0}$), which may be attributed to periodically varying $\kappa(t)$ even at long times. We present a dynamical phase diagram w.r.t. $\epsilon$ and $m$ (Cf. Fig. 3). The transition from Markovian to non-Markovian regime at $m=2\kappa_{0}$ is marked by TISS for all values of $\epsilon$ in the Markovian regime for large enough $\kappa_{0}$, which changes to varied dynamical phases, including DTC, in the non-Markovian regime. Furthermore, the DTC phase emerging for $m$ close to $2\kappa_{0}$ shows substantial robustness w.r.t. $\epsilon$, while this robustness is reduced, even though does not vanish, as we move deeper into the non-Markovian regime of $m<2\kappa_{0}$. Furthermore, we verify the presence of robust DTC order in the non-Markovian regime even for aperiodic $\kappa(t)$, realized by introducing a random noise in the dissipator coefficient. Our analysis involves different time scales, which we tabulate below in Table 1.
@@
We expect the dissipative DTCs studied here can be realized experimentally in currently existing setups. For example, the presence of DTCs has been verified experimentally in different platforms, both in closed quantum systems, such as in ion traps [8], and in the presence of dissipation, for example in optical cavities [45, 46]. Experimental studies of DTCs in the presence of different forms of dissipation would require control over bath spectral functions, which for example can be achieved through the introduction of filters [47].
Table 1: Time scales
Time scales
Markovian (M)
regime
Non Markovian(NM)
regime
Remarks
Modulation period of $\lambda_{t}$
$T$ $T$
In the DTC phase the system shows oscillations with
a time period $2T$
Thermalization time $\tau_{\rm th}$
$\tau_{\rm th}\sim\kappa_{0}^{-1}$ $\tau_{\rm th}\sim\kappa_{0}^{-1}$
DTC phase is preserved for intermediate values of $\kappa_{0}$
($\kappa_{0}\gtrsim T^{-1}$) in the NM regime.
Time period of information
back flow $\equiv$ time period of $\kappa(t)$
NA $T_{NM}=4\pi/|d|$
In the NM regime we have taken $T_{\rm NM}=T$ for simplicity
Our results show that non-Markovian dynamics can be highly relevant for time-translational symmetry breaking in many-body open quantum systems, and can show features distinct from those seen in the Markovian regime. This also raises open questions regarding the fate of DTCs in more generic forms of non-Markovian dynamics, and regarding the possible role played by information back-flow in the formation of DTCs.
@@ Appendix A Evolution equations and steady states
Appendix B Dynamics for large $\kappa_{\rm{max}}$
In case of large $\kappa_{\rm{max}}$ dissipation dominates the dynamics, such that the system reaches a TISS, as shown in Fig. 5 a) or a thermal phase, as shown in Fig. 5 b).
Figure 5: Stroboscopic dynamics for $\kappa_{\rm{max}}=10$, $\epsilon=0.02$ , a) $m=\kappa_{0}$ and b) $m=\frac{\kappa_{0}}{4}$.
Appendix C Limit cycle
As we vary $\epsilon$ the system may oscillate between two limit cycles, as shown in Fig. 6 (see Sec. II.2.2).
Figure 6: A limit cycle, Stroboscopic dynamics for $\kappa_{\rm{max}}=3$, $\epsilon=0.07$ , a) $m=\frac{\kappa_{0}}{5}$.